Circular No. 7616

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 A2 (LINEAR) C. W. Hergenrother, M. Chamberlain, and Y. Chamberlain, Lunar

and Planetary Laboratory, University of Arizona, report that 60-s R-band images of C/2001 A2 taken on Apr. 30.12 UT with the Catalina 1.54-m reflector show a double nucleus. The two components are nearly equal in brightness and 3′′.5 apart and aligned precisely on an east-west line. Both components are highly condensed. Observations (with the same telescope) on Apr. 24.14 show only a single nucleus.

Visual m_{1} estimates: Apr. 20.93 UT, 7.2 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars); 21.46, 7.4 (M. Mattiazzo, Wallaroo, S. Australia, 7×50 binoculars); 23.16, 6.7 (C. S. Morris, Fillmore, CA, 20×80 binoculars); 25.35, 6.4 (S. T. Rae, Whakatane, New Zealand, 8×21 binoculars); 27.48, 6.4 (A. Pearce, Nedlands, W. Australia, 8×40 binoculars); 29.95, 6.3 (W. Souza, São Paulo, Brazil, 11×80 binoculars); 30.49, 6.3 (Pearce).

SUPERNOVA 2001bb IN IC 4319 R. Chornock, M. Modjaz, and A. V. Filippenko, University of Cal-

ifornia at Berkeley, write: “Inspection of a noisy CCD spectrum (range 320–1000 nm), obtained on Apr. 30 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2001bb (cf. IAUC 7614) is a supernova, probably of type Ic, but perhaps type Ia. No obvious H or He lines are visible, although weak Si II 615-nm absorption might be present. The classification is quite uncertain at present.”

P. Nugent and L. Wang, Lawrence Berkeley National Laboratory, report: “A spectrum (range 400–800 nm) obtained on Apr. 30.21 at the Kitt Peak 2.1-m telescope (+ GCAM) shows that SN 2001bb is a type-I supernova. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4653 km/s for the host galaxy, the minimum of the Si II feature (rest 635.5 nm) is at a velocity of 13 800 km/s. Ca II, O I, and Fe II features also appear at this velocity. Several interesting absorption features between 500 and 600 nm have yet to be identified conclusively, leaving its subtype unknown at present. The supernova most closely resembles SN 1999as (cf. IAUC 7128), a type-Ic hypernova event (cf. Paczyński 1998, Ap.J. 494, L45) that peaked at absolute V magnitude brighter than −21.”

2001 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green